OpenStudy (anonymous):

The equation for an ellipse is \[\frac{ x ^{2} }{ a ^{2} }+\frac{ y ^{2} }{ b ^{2}}=1\] and the foci is given by (c,0) where \[c ^{2}=a ^{2}-b ^{2}\] and my question is if b>a is the foci given by (0,c) where \[c ^{2}=a ^{2}-b ^{2}\]

OpenStudy (asnaseer):

yes, if a>b then foci are at \((\pm c, 0)\) otherwise they are at \((0, \pm c)\). see here for more details: http://hotmath.com/hotmath_help/topics/ellipse.html

OpenStudy (anonymous):

I suppose you could try to derive the equation as a locus from the two foci (c,0) and (-c,0). \[\sqrt{y^2+(x-c)^2}+\sqrt{y^2+(x+c)^2}=k\]

OpenStudy (anonymous):

It seems you're confused with the notation: Foci always belong to major axis, and major axis always named a!

OpenStudy (anonymous):

\[y^2+(x-c)^2+y^2+(x+c)^2+\sqrt{y^2+(x-c)^2}\sqrt{y^2+(x+c)^2}=k^2\] \[y^2+(x-c)^2+y^2+(x+c)^2-k^2=-\sqrt{y^2+(x-c)^2}\sqrt{y^2+(x+c)^2}\] \[(y^2+(x-c)^2+y^2+(x+c)^2-k^2)^2=(y^2+(x-c)^2)(y^2+(x+c)^2)\]ans so on

OpenStudy (anonymous):

*and

OpenStudy (anonymous):